OpenStudy (anonymous):

my trajectory is y = 3x + 6 ----- (1) The IH orbit is: y^2 + x^2 = 40,000 ------- (2) Solve these two equations to find the points where the two paths intersect.

OpenStudy (anonymous):

here is the assignment

OpenStudy (mertsj):

\[(3x+6)^2+x^2=40000\] \[9x^2+36x+36+x^2=40000\] \[10x^2+36x+36-40000=0\] \[10x^2+36x-39964=0\]

OpenStudy (mertsj):

\[5x^2+18x-19982=0\]

OpenStudy (anonymous):

a = 1, b = 3.6, c = -3996.4 ?

OpenStudy (anonymous):

we plug that into the quadratic equation right

OpenStudy (mertsj):

\[x=\frac{-9\pm \sqrt{99991}}{5}\]

OpenStudy (anonymous):

x=-65 or 61.4

OpenStudy (anonymous):

correct

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

and for the y values its -189 and 190.2 ?

OpenStudy (anonymous):

^^^

OpenStudy (mertsj):

\[y=\frac{-27\pm \sqrt{99991}+30}{5}=\frac{3\pm \sqrt{99991}}{5}\]

OpenStudy (anonymous):

? are the y values correct

OpenStudy (anonymous):

hm ok so the lines intersect at (-65,63.8)

OpenStudy (anonymous):

-189 and 190.2 lol

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

Those are right.

OpenStudy (anonymous):

ok,so the point is (-65, -189)

OpenStudy (mertsj):

And (61,190)

OpenStudy (mertsj):

The line intersects that circle in two points.

OpenStudy (anonymous):

ohh i see ok

OpenStudy (anonymous):

so that is for #3 correct ?

OpenStudy (mertsj):

I don't know. You said find the intersection of those two equations and that's what we did.

OpenStudy (anonymous):

ok thank you . you are a great help!

OpenStudy (mertsj):

Yes that should be number 3

OpenStudy (anonymous):

:) thanks man

OpenStudy (anonymous):

or woman lol

OpenStudy (mertsj):

yw